# Do Robots Really Need Anthropomorphic Hands?

### v1

@@ 2.3.6 Sensorimotor Learning
[116] distinguish various processes of motor learning: (1) error-based learning that uses a gradient with respect to each component of the motor command to adapt motor commands (e.g., grip force) quickly, (2) reinforcement learning for complex sequences of actions that need to take place to achieve a goal and the outcome is far removed from the action, and (3) use-dependent learning that changes the behavior, even if no outcome information is available. In addition, humans can learn high-level information about which movements to make in sequence and how to compensate for perturbations from observing others.
@@ 2.4.2 Perception
The human hand robustly integrates tactile and proprioceptive signals across skin, muscle, and joint receptors. This integration enables the perception of detailed object properties and nuanced positional feedback, forming a sophisticated multimodal sensory system. Such complexity in perception is a key inspiration for developing advanced robotic hands that mimic human dexterity and precision.
@@ 2.4.3 Control and Learning
Different mechanisms seem to be responsible for controlling and acquiring fast, complex, repetitive skills without sensor input, versus new or perception-based skills that require constant control. Often, human control aims at optimality and does not explicitly encode which limb or muscle is used. Humans integrate sensor data from various high-dimensional sources (visual, tactile, kinesthetic) and easily coordinate many degrees of freedom. They can handle sensor and actuator noise in a Bayesian-like way and anticipate sensory feedback. However, the most important feature of human control is its adaptability through learning and continuous improvement with additional experience.
@@ 3.1 Mechanics
Robotic hands are designed in various ways depending on their intended use. [16] suggest considering six aspects when designing a robotic hand: (1) hand kinematics, (2) actuation principle, (3) actuation transmission, (4) sensors, (5) materials, and (6) manufacturing method.
@@
[16] provides six possible kinematic architectures in robotic hands, depicted in Figure 6.
Figure 6: Representation of possible robotic hand kinematic architectures, adapted from [16]. $M$ denotes the number of actuated tendons (red circles), and $N$ the number of bidirectional DoFs (black circles). Diagrams show: (a) $M=N$ coupled joints; (b) $M<N$ underactuated transmission; (c) $M=N$ fully actuated open chain; (d) $M=N$ fully actuated closed chain; (e) $M=N+1$ fully controllable; (f) $M=2N$ agonist/antagonist transmission [16]. Continuous lines indicate tendons, dashed lines joints, and zigzag lines springs.
The Shadow Dexterous Hand is one of the most advanced robotic hands and is produced by Shadow Robot. It has five fingers and 24 DoFs, of which 20 are controllable. In addition, similar to the human hand, it is possible to perform opposition and reposition of the thumb, flexion, extension, adduction, and abduction of all fingers. Such movements enable a dexterous hand that can perform various complex manipulation tasks.
Figure 7(c) represents the kinematic architecture of the Shadow Dexterous Hand using the kinematic notations derived from [110] and depicted in Figure 7(a).
@@
Figure 7: Kinematic diagrams. Notation based on [110].
@@
Apart from the chosen kinematic architecture, the transmission actuator mechanism and kind do impact the performance of the robotic hand. For instance, actuators used in the artificial hand can be electrical (e.g., DC motors), pneumatic, or hydraulic. Natural muscles have a power density ($\rho$) of approximately $500W/Kg$. Such power density can be achieved by hydraulic or pneumatic actuators (respectively, $2000W/Kg$ and 400W/Kg400W/Kg[32]. Such solutions are already used in artificial robotic hands (e.g., GRIPKIT Industrial (P PRO) [^4] https://weiss-robotics.com/gripkit-for-cobots/, EHA hand [45]). However, such solutions are usually bulky or hard to maintain. DC motors (electrical actuators), on the other hand, can reach a power density $\rho=100W/Kg$ [32]. While these actuation mechanisms may have a lower power-to-weight ratio compared to others, they remain popular actuators thanks to their compact size and cost-effective maintenance [16].
In order to accurately control joints, transmission is crucial. Such systems enable converting the power provided by the actuators to a specific hand/finger movement. An ideal transmission system would have low inertia, friction and backlash, and would be overall compact and small (i.e., in terms of size and weight). Several transmission types can be deployed: including tendons, gear trains, belts, linkages or flexible shafts [16]. In the human hand, we found tendons that connect muscles to bones [98]. Such tendons run in a sheath. A similar system is commonly found in artificial robotic hands, as such a design allows for reduced friction and remote control of the fingers (e.g., Shadow dexterous hand). For further transmission comparison, we refer to the work of [16].
@@
We focus on six different features: DoF, number of actuators, number of fingers, possibility of flexion (and extension), adduction (and abduction), and opposition (and reposition). Specifically, DoF and the number of actuators influence the hand’s dexterity and versatility, while the number of fingers and their movement capabilities affect its grasping and manipulation abilities. For a detailed description of each robotic hand, please refer to the supplementary material. The characteristics of robotic and prosthetic hands are visualized in Figure 8.
@@ 3.2 Perception
To perform dexterous manipulation, a robot must understand what the object being manipulated is and what type of operation is required, i.e., task requirements [117].
Tactile sensors are primarily designed to mimic mechanoreceptors, particularly for detecting mechanical pressure. The main objectives of tactile sensors are to determine the location, shape, and intensity of contacts. These properties are determined by measuring the instantaneous pressure or force applied to the sensor’s surface on multiple contact points. Also, the contact’s late effects, i.e., body-borne vibrations, may carry relevant information. Body-borne vibrations are not as commonly measured or exploited as part of haptic sensing; however, there are some examples and it is becoming an active area of research [e.g., 37, 9, 107], including sensors that are inspired by the hair follicle receptors or ciliary structure [3, 39] and that have been proven effective in obtaining information about the texture of objects [80, 79].
@@
Overall, artificial tactile sensors are much less established than artificial visual receptors, e.g., RGB(-D) cameras, and event-based cameras. Technologies for tactile sensing have been developed since the early 1970s and have undergone significant improvements in the past decade [18, 17, 42]. However, the field remains young, and there are no widely accepted solutions. Several transduction methods have been explored, including capacitive [51], piezoelectric [90], piezoresistive [38], optical [112, 47], fiber optics [77], and magnetic [34]. Table 1 summarizes the advantages and disadvantages of the different transduction principles for detecting mechanical pressure. For additional information, please refer to [13].
@@ 3.2.1 Commercial Sensors
Seed Robotics’ FTS Tactile pressure sensors (see Figure 9) are low-cost sensors with high-resolution contact force measurement capabilities (1mN to 30N range). The sensor compensates for temperature and is immune to magnetic interference.
@@
Figure 9: Left: the SINGLEX stand-alone tactile pressure sensor version. Right: FTS tactile pressure sensor mounted on a robot finger. Images used with permission from Seed Robotics (https://www.seedrobotics.com/).
The uSkin sensor [104] by Xela Robotics is a magnetic tactile sensor composed of small magnets embedded in a thin layer of flexible rubber and placed above a matrix of magnetic Hall-effect sensor chips. Upon contact, the magnets are displaced, and the magnetic field sensed by the Hall-effect chips changes; the contact forces can be estimated from these variations in the magnetic field. The uSkin sensor can measure the full 3D force vector (i.e., both normal and shear contact forces) at each tactel, with a good spatial resolution (about 1.6 tactels for square cm), high sensitivity (minimum detectable force of 1gf), and high frequency (>100>100Hz, depending on the configuration). Different versions of the sensor are available to cover both flat and multi-curved surfaces, see Figure 10 for an example.
@@
Figure 10: Left: a flat version inspired by [105]. Right: a multi-curved version inspired by [106]. Images with permission from Xela Robotics (https://xelarobotics.com/).
The GelSight Mini [119] and DIGIT [49] tactile sensors by GelSight are optical tactile sensors that utilize a piece of elastomeric gel with a reflective membrane coating on top, enabling them to capture fine geometrical textures as deformations in the gel. A series of LEDs with RGB colour illuminates the gel, allowing a camera to record the deformation.
Finally, Contactile offers both a stand-alone sensor and tactile sensor arrays, called PapillArray sensors, as shown in Figure 11. These optical sensors consist of infrared LEDs, a diffuser, and four photodiodes encapsulated in a soft silicone membrane. The photodiodes are used to measure the light intensity patterns, which are then used to infer the displacement and force applied to the membrane. This strategy enables the measurement of 3D deflections, 3D forces, and 3D vibrations, as well as the inference of emergent properties such as torque, incipient slip, and friction.
@@
Figure 11: Left: Single 3D force tactile sensor. Right: A slim tactile sensor array (PapillArray Sensor) available in different configurations. Images from Contactile (https://contactile.com/), licensed under $C$ BY-NC-ND 4.0.
@@
The need for such technologies is driving research forward in both the development of new sensing technologies and applications such as robotic grasping, smart prostheses, and surgical robots. In particular, enhancements are still needed in the mechanical robustness, sensitivity, and reliability of the measurements, as well as the ease of electromechanical integration and replacement, to deploy sensors in practical applications. Of particular interest are solutions that: are flexible [51, 91], stretchable [11] and can cover sizeable [19] and curved [106] surfaces (possibly with a small number of electrical connections [37]), can detect multiple contacts at the same time [31], can detect both normal and shear forces [105], are affordable and can be easily manufactured [74]. For more information on experimental tactile sensing technologies, see [13], and for a specialized review of printable, flexible, and stretchable tactile sensors, see [91].
@@ 3.2.2 Tactile Servoing
In the process of dexterous manipulation, the tactile servo can be assigned in the control architecture, where the motor is driven by tactile feedback. Meanwhile, tactile servoing also gives the robotic hand a new direction for exploring new environments more safely [117].
@@
Dexterous manipulation with a multi-fingered robotic hand is a challenging task due to the existence of uncertainties arising from sensor noise, slippage, or external disturbances. External disturbances caused by environmental changes may cause a planned-to-be stable grasp to become unstable. Humans can react quickly to instabilities through tactile sensing. The ability to detect slips in tactile perception is critically essential for further dexterous manipulations [117].
@@ 3.3 Robot Control
Numerous approaches have been used to generate motor skills for robots. A list that is most likely not complete: motion and grasp planning, optimal control, model-predictive control, reinforcement learning (model-free / model-based), inverse reinforcement learning, offline reinforcement learning, imitation learning (e.g., through teleoperation, kinesthetic teaching, and retargeting), and evolutionary algorithms. Robotics research and neuroscience have a significant overlap as demonstrated, for example, by [94] or [101]. Theories that explain human behavior often yield practically useful concepts for movement generation in robotics.

### v2

@@ 2.3.6 Sensorimotor Learning
Wolpert et al. [48] distinguish various processes of motor learning: (1) error-based learning that uses a gradient with respect to each component of the motor command to adapt motor commands (e.g., grip force) quickly, (2) reinforcement learning for complex sequences of actions that need to take place to achieve a goal and the outcome is far removed from the action, and (3) use-dependent learning that changes the behavior, even if no outcome information is available. In addition, humans can learn high-level information about which movements to make in sequence and how to compensate for perturbations from observing others.
@@ 2.4.2 Perception
The human hand robustly integrates tactile and proprioceptive signals across skin, muscle, and joint receptors. This integration enables the perception of detailed object properties and nuanced positional feedback, forming a sophisticated multimodal sensory system. Moreover, precise measurement of contact forces is essential for force control and for maintaining stable grasps [17]. The Lack of proprioception can lead to unstable reach trajectories, impaired postural stability, and poor hand coordination, even when visual feedback is available [19].
@@ 2.4.3 Control and Learning
Different mechanisms seem to be responsible for controlling and acquiring fast, complex, repetitive skills without sensor input, versus new or perception-based skills that require constant control. Often, human control aims at optimality (in skill performance) and does not explicitly encode which limb or muscle is used. Humans integrate sensor data from various high-dimensional sources (visual, tactile, kinesthetic) and easily coordinate many degrees of freedom. They can handle sensor and actuator noise in a Bayesian-like way and anticipate sensory feedback. However, the most important feature of human control is its adaptability through learning and continuous improvement with additional experience.
@@ 3 Robotic Hands
Figure 6: Representation of possible robotic hand kinematic architectures, adapted from Controzzi et al. [8]. $M$ denotes the number of actuated tendons (red circles), and $N$ the number of bidirectional DoFs (black circles). Diagrams show: (a) $M=N$ coupled joints; (b) $M<N$ underactuated transmission; (c) $M=N$ fully actuated open chain; (d) $M=N$ fully actuated closed chain; (e) $M=N+1$ fully controllable; (f) $M=2N$ agonist/antagonist transmission [8]. Continuous lines indicate tendons, dashed lines joints, and zigzag lines springs.
@@ 3.1 Mechanics
Robotic hands are designed in various ways depending on their intended use. Controzzi et al. [8] suggest considering six aspects when designing a robotic hand: (1) hand kinematics, (2) actuation principle, (3) actuation transmission, (4) sensors, (5) materials, and (6) manufacturing method.
@@
Controzzi et al. [8] provides six possible kinematic architectures in robotic hands, depicted in Figure 6.
The Shadow Dexterous Hand is one of the most advanced robotic hands and is produced by Shadow Robot. It has five fingers and 24 DoFs, of which 20 are controllable. In addition, similar to the human hand, it is possible to perform opposition and reposition of the thumb, flexion, extension, adduction, and abduction of all fingers. Such movements enable a dexterous hand that can perform various complex manipulation tasks. Figure 7(c) represents the kinematic architecture of the Shadow Dexterous Hand using the kinematic notations derived from [49] and depicted in Figure 7(a).
@@
Figure 7: Kinematic diagrams. Notation based on Wang at el. [49].
@@
A central challenge in hand design is the optimal number of fingers required for stable grasping versus in-hand manipulation. While five fingers (as seen in the Shadow Dexterous Hand) maximize anthropomorphism, they introduce significant control complexity and mechanical packaging challenges. Conversely, three-fingered hands (like the BarrettHand) are frequently employed to achieve stable tripod grasps, offering a significant dexterity improvement over simple parallel-jaw grippers. However, non-anthropomorphic architectures with four fingers were also considered: designs such as the Yale OpenHand [50] or the Tesollo DG-4F hand [51] introduce the mechanical redundancy necessary for finger gaiting, by using two opposing finger pairs. This configuration reduces the actuator count compared to a five-fingered hand, yet provides significantly higher manipulability and grasp safety than a three-fingered gripper, particularly during object reorientation tasks. The non-anthropomorphic concept of two opposing fingers mitigates obstructions and workspace reduction during in-hand manipulation.
Apart from the chosen kinematic architecture, the transmission actuator type and mechanism do impact the performance of the robotic hand. For instance, actuators used in the artificial hand can be electrical (e.g., DC motors), pneumatic, or hydraulic. Natural muscles have a power density ($\rho$) of approximately $500W/Kg$. Such power density can be achieved by hydraulic or pneumatic actuators (respectively, $2000W/Kg$ and 400W/Kg400W/Kg[52]. Such solutions are already used in artificial robotic hands (e.g., GRIPKIT Industrial (P PRO) [^7] https://weiss-robotics.com/gripkit-for-cobots/, EHA hand [53]). However, such solutions are usually bulky or hard to maintain. DC motors (electrical actuators), on the other hand, can reach a power density $\rho=100W/Kg$ [52]. While these actuation mechanisms may have a lower power-to-weight ratio compared to others, they remain popular actuators thanks to their compact size and cost-effective maintenance [8].
In order to accurately control joints, transmission is crucial. Such systems enable converting the power provided by the actuators to a specific hand/finger movement. An ideal transmission system would have low inertia, friction and backlash, and would be overall compact and small (i.e., in terms of size and weight). Several transmission types can be deployed: including tendons, gear trains, belts, linkages or flexible shafts [8]. In the human hand, we found tendons that connect muscles to bones [54]. Such tendons run in a sheath. A similar system is commonly found in artificial robotic hands, as such a design allows for reduced friction and remote control of the fingers (e.g., Shadow dexterous hand). For further transmission comparison, we refer to the work of Controzzi et al. [8].
@@
We focus on six different features: DoF, number of actuators, number of fingers, possibility of flexion/extension, adduction/abduction, and opposition/reposition. Specifically, DoF and the number of actuators influence the hand’s dexterity and versatility, while the number of fingers and their movement capabilities affect its grasping and manipulation abilities. For a detailed description of each robotic hand, please refer to the supplementary material. The characteristics of robotic and prosthetic hands are visualized in Figure 8.
@@
To synthesize the design choices discussed above, Table 1 summarizes the key trade-offs between different kinematic and actuation architectures found in current state-of-the-art robotic and prosthetic hands. Furthermore, robotic hands are attached to robotic arms. Six joints are sufficient to control all position and orientation components in 3D space so that more DOF are redundant. However, a seventh, redundant DOF allows for avoiding singularities, obstacles, joint limits, or torque limits [55, 56]. Hence, arm design should not be neglected when developing dexterous manipulation systems.
Table 1: Comparison of common robotic hand design architectures: Advantages and Disadvantages.
Design Architecture Advantages Disadvantages
Actuation Principle
Fully Actuated (e.g., Barrett Hand) High dexterity and manipulability Independent control of each joint Capable of complex in-hand manipulation High control complexity (high dimensionality) Increased weight, cost, and size Higher energy consumption
Underactuated (e.g., Prosthetic) Mechanical simplicity and lower weight Passive adaptability to object shape Lower cost and easier control Reduced dexterity (cannot control all joints) Difficulty performing precise in-hand manipulation Predictability of grasp can be lower
Transmission Type
Tendon-Driven Motors can be placed remotely (forearm) reducing hand inertia Mimics biological systems Allows for slim finger design Issues with friction, elasticity, and hysteresis Complex maintenance (tendon routing/breakage) Requires tensioning mechanisms
Linkage/Gear-Driven High structural stiffness and precision Robust force transmission Less maintenance required than tendons Higher backlash (unless direct drive) Actuators usually located in the hand, increasing inertia Rigid collisions can damage gears
Beyond the established categories and traditional kinematic designs, the state of the art is currently shifting toward three major hardware development trends driven by computational and mechanical necessities:
  • Soft Robotics: There is a distinct move toward compliant materials and fluidic or pneumatic actuation. By utilizing elastomers and variable stiffness mechanisms, these hands ensure safe human-robot interaction and passive adaptability to object geometries, reducing control complexity for delicate grasping tasks [57, 58].
  • AI-Driven Hardware Design: The integration of Artificial Intelligence (AI) and Reinforcement Learning (RL) is heavily influencing hardware specifications. Modern hands are increasingly designed to be mechanically robust to withstand the high-cycle, trial-and-error collisions inherent to learning algorithms. Furthermore, “hardware lottery” concepts are driving the co-design of morphology and policy, ensuring that the physical hand simplifies the learning landscape for the agent [59].
  • Direct and Quasi-Direct Drive Actuation: To prioritize bandwidth and back-drivability, recent designs utilize high-torque motors with either no gearing (Direct Drive) or low gear ratios (Quasi-Direct Drive, typically <10:1<10:1). This architecture provides impact transparency, allowing for force estimation without expensive torque sensors. Notable examples include the CMU Direct Drive Hand [60], which uses a 1:1 transmission for maximum transparency, and the Blue grippers [61], which utilize QDD to balance torque and compliance.
@@ 3.2 Perception
To perform dexterous manipulation, a robot must understand what the object being manipulated is and the type of operation required, i.e., the task requirements [62].
Tactile sensors are primarily designed to mimic mechanoreceptors, particularly for detecting mechanical pressure. The main objectives of tactile sensors are to determine the location, shape, and intensity of contacts. These properties are determined by measuring the instantaneous pressure or force applied to the sensor’s surface on multiple contact points. Also, the contact’s late effects, i.e., body-borne vibrations, may carry relevant information. Body-borne vibrations are not as commonly measured or exploited as part of haptic sensing; however, there are some examples and it is becoming an active area of research [e.g., 63, 64, 65, 66], including sensors that are inspired by the hair follicle receptors or ciliary structure [67, 68] and that have been proven effective in obtaining information about the texture of objects [69, 70].
@@
Overall, artificial tactile sensors are much less established than artificial visual receptors, e.g., RGB(-D) cameras, and event-based cameras. Technologies for tactile sensing have been developed since the early 1970s and have undergone significant improvements in the past decade [17, 20, 75]. However, the field remains young, with no widely accepted technical solutions and no unified tactile representation yet. Table 2 summarizes the advantages and disadvantages of the different transduction principles for detecting mechanical pressure. For additional information, please refer to [76].
@@ 3.2.1 Commercial Sensors
Seed Robotics’ FTS Tactile pressure sensors (see Figure 9(a)) are low-cost sensors with high-resolution contact force measurement (1mN to 30N) and a sample frequency of 50Hz. The sensor compensates for temperature and is immune to magnetic interference.
@@
(a) Left: the SINGLEX stand-alone tactile pressure sensor version. Right: FTS tactile pressure sensor mounted on a robot finger. Images used with permission from Seed Robotics (https://www.seedrobotics.com/).
@@
(b) Left: a flat version inspired by [83]. Right: a multi-curved version inspired by [84]. Images with permission from Xela Robotics (https://xelarobotics.com/).
@@
(c) Left: Single 3D force tactile sensor. Right: A slim tactile sensor array (PapillArray Sensor) available in different configurations. Images from Contactile (https://contactile.com/), licensed under $C$ BY-NC-ND 4.0.
Figure 9: Tactile sensors.
The uSkin sensor [85] by Xela Robotics is a magnetic tactile sensor composed of small magnets embedded in a thin layer of flexible rubber placed above a matrix of magnetic Hall-effect sensor chips. Upon contact, the magnets are displaced, and the magnetic field sensed by the Hall-effect chips changes; the contact forces can be estimated from these changes. The uSkin sensor can measure the full 3D force vector (i.e., both normal and shear contact forces) at each tactel, with a good spatial resolution (about 1.6 tactels for square cm), high sensitivity (minimum detectable force of 1gf), and high frequency (>100>100Hz, depending on the configuration). Different versions of the sensor are available to cover both flat and multi-curved surfaces, see Figure 9(b) for an example.
The GelSight Mini [86] and DIGIT [87] tactile sensors by GelSight are optical tactile sensors that utilize a piece of elastomeric gel with a reflective membrane coating on top, enabling them to capture fine geometrical textures as deformations in the gel. A series of LEDs with RGB colour illuminates the gel, allowing a camera to record the deformation and output it at up to 30Hz.
Finally, Contactile offers both a stand-alone sensor and tactile sensor arrays, called PapillArray sensors, as shown in Figure 9(c). These optical sensors consist of infrared LEDs, a diffuser, and four photodiodes encapsulated in a soft silicone membrane. The photodiodes measure light-intensity patterns, which are then used to infer the membrane’s displacement and force applied to it. This strategy enables the measurement of 3D deflections, 3D forces, and 3D vibrations, as well as the inference of emergent properties such as torque, incipient slip, and friction. The sensor can measure 3D applied forces up to 20N at a rate of up to 1kHz.
@@
The need for such technologies is driving research forward in both the development of new sensing technologies and applications such as robotic grasping, smart prostheses, and surgical robots. In particular, enhancements are still needed in the mechanical robustness, sensitivity, and reliability of the measurements, as well as the ease of electromechanical integration and replacement, to deploy sensors in practical applications. Of particular interest are solutions that: are flexible [77, 88], stretchable [89] and can cover sizeable [90] and curved [84] surfaces (possibly with a small number of electrical connections [63, 64]), can detect multiple contacts at the same time [91], can detect both normal and shear forces [83], are affordable and can be easily manufactured [92]. For more information on experimental tactile sensing technologies, see [76], and for a specialized review of printable, flexible, and stretchable tactile sensors, see [88].
@@ 3.2.2 Tactile Servoing
In the process of dexterous manipulation, the tactile servo can be assigned in the control architecture, where the motor is driven by tactile feedback. Meanwhile, tactile servoing also provides the robotic hand with a new direction for exploring environments more safely [62].
@@
Dexterous manipulation with a multi-fingered robotic hand is a challenging task due to the existence of uncertainties arising from sensor noise, slippage, or external disturbances. External disturbances, such as environmental changes, may cause a planned-to-be stable grasp to become unstable. Humans can react quickly to instabilities through tactile sensing. The ability to detect slips in tactile perception is critically essential for further dexterous manipulations [62].
@@ 3.3 Robot Control
Since our focus in this article is on mechanics and perception, we only briefly summarize the most relevant subfields of robot control and specifically focus on examples of control for dexterous manipulation. Numerous approaches have been used to generate motor skills for robots. A list that is most likely not complete: motion and grasp planning, optimal control, model-predictive control, reinforcement learning (model-free / model-based), inverse reinforcement learning, offline reinforcement learning, imitation learning (e.g., through teleoperation, kinesthetic teaching, and retargeting), and evolutionary algorithms. Robotics research and neuroscience have a significant overlap as demonstrated, for example, by Sternad [35] or Todorov [41]. Theories that explain human behavior often yield practically useful concepts for movement generation in robotics.